Find the exact value of x?
arcsinx - arccosx = pi/6
our prof said in this case we should let theta=arccosx but after that everything is all blurry and I just don't know how to go about this. Someone please explain to me how to get the answer. Also please show the steps if you can. Thank you!

Answer by Alan3354(69443)

Find the exact value of x?
arcsin(x) - arccos(x) = pi/6
arcsin(x) + arccos(x) = pi/2 Identity
------------------------------ Add
2arcsin(x) = 2pi/3
arcsin(x) = pi/3
x = sqrt(3)/2